OpenStudy (countrygirl1431):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (countrygirl1431):

@rishavraj

rishavraj (rishavraj):

u got P and Q.....so find first of all find the slope..

OpenStudy (countrygirl1431):

thats y=mx+b right

rishavraj (rishavraj):

\[\frac{ -4 - 2 }{ 8 - 6 } = m\]

rishavraj (rishavraj):

now just use \[\frac{ y - 2 }{ x - 2 } = m \]

OpenStudy (countrygirl1431):

i still dont get it